# Protecting Quantum Information via Many-Body Dynamical Localization

### v2

@@ Abstract
Dynamically localized states in quantum many-body systems are fundamentally important in understanding quantum thermalization and have applications in quantum information processing. Here we explore many-body dynamical localization (MBDL) without disorders in a non-integrable quantum XY spin chain under periodical and quadratic kicks. We obtain the localization phase diagram with the MBDL and delocalization states and show dynamical observables to extract the phase diagram. For proper kick strengths in the MBDL regime, we reveal a local dynamical decoupling effect for persistent Rabi oscillation of certain spins. Furthermore, we propose the MBDL-protected quantum information at high temperatures, and present an analysis of the dynamical decoupling to obtain the required system parameters for quantum storage. Compared to other non-thermalized states, the disorder-free MBDL states require much fewer repetitions and resources, providing a promising way to protect and store quantum information robust against thermal noises.
Introduction.— Coherent and controllable many-body quantum dynamics plays an important role in quantum matter engineering and quantum information processing. As a consequence of thermalization, however, coherence in generic many-body quantum systems is fragile once they are excited to a finite energy density above ground states Rigol et al. (2008); D’Alessio et al. (2016). There are several ways to avoid thermalization, apart from integrability. One is the many-body localized (MBL) states induced by sufficiently strong disorders Anderson (1958); Nandkishore and Huse (2015); Abanin et al. (2019); Sierant et al. , and the other is quantum many-body scars Bernien et al. (2017); Turner et al. (2018); Bluvstein et al. (2021); Serbyn et al. (2021); Zhang et al. (2022a), which correspond to the strong and weak breaking of ergodicity, respectively. In MBL systems, all excitations are localized and thermalization is strongly suppressed, which can protect coherence of any initial states. The scar states also have a long coherence time as their dynamics are effectively restricted in a subspace. These localized states may provide resources for quantum information processing Huse et al. (2013); Serbyn et al. (2013a, 2014); Bahri et al. (2015); Yao et al. ; Zhang et al. (2018); Dong et al. (a); Naik et al. (2019); Mi and et al. (2022); Schmid et al. (2024); Yarloo et al. (2024).
@@
In this work, we investigate the disorder-free MBDL in a non-integrable kicked XY spin chain and its application in protecting quantum information. For the quantum XY chain with a longitudinal field under periodic transverse-field kicks, we obtain the phase diagram with the MBDL and delocalization phases and show dynamical observables to extract the localization phase diagram. We uncover a local dynamical decoupling (DD) effect under proper kick strengths in the MBDL regime, which enables persistent Rabi oscillation of certain spins. Moreover, we demonstrate the MBDL-protected quantum information at high temperatures, and present an analysis of the local DD to obtain the system parameters for robust quantum storage. Our proposal is realizable with superconducting and Rydberg atom quantum simulators.
@@
where $H_{\text{XY}}=\sum_{j=1}^{L}J\left(\sigma_{j}^{+}\sigma_{j+1}^{-}+\sigma_{j+1}^{+}\sigma_{j}^{-}\right)+\Omega\sigma_{j}^{x}$ is the static Hamiltonian with XY interaction between nearest-neighbour spins and a longitudinal field, $K$ and $T$ denote the strength and period of the transverse-field kicks with a quadratic potential centered at $j_{0}$, respectively. Here $\sigma_{j}^{\pm}=\sigma_{j}^{x}\pm i\sigma_{j}^{y}$ with $\sigma_{j}^{x,y,z}$ being the Pauli matrices of $j$-th spin, $J$ is the interaction strength, and $\Omega$ is the strength of the longitudinal field, which leads to coherent spin flips and breaks $U(1)$ symmetry. Notably, we will emphasize in protection of quantum coherence raising from $\Omega$ via the MBDL in this kicked XY chain, which was not considered in kicked Ising chains Prosen (2002); Bertini et al. (2018, 2019); Khasseh et al. (2023); Sharma and Bhosale (2024). Hereafter we consider the high-frequency driving by setting $T=1/16$ and $\Omega=\hbar=1$ as the energy (frequency) unit.
@@
where $H_{F}=\frac{i}{T}\ln U_{F}$ is the effective Floquet Hamiltonian. The wave function at time $t=nT$ after $n$ kick cycles is |$\ket{\psi(t)}=U_{F}^{n}\ket{\psi(0)}$. As the model is non-integrable and $U(1)$-symmetry broken, we study the static and dynamical properties using the exact diagonalization Weinberg and Bukov (2019) for system size $L\leq 14$ and density matrix renormalization group White (1992); Schollwöck (2011); Fishman et al. (2022) for larger $L$. Without loss of generality, we consider centrosymmetric spin chains by setting $j_{0}=(L+1)/2$ under open boundary condition.
Localization phase diagram.— We first study localization properties of many-body Floquet eigenstates {|$\{\ket{\psi^{F}_{\alpha}}\}$. They can be numerically obtained from the eigen equation $U_{F}\ket{\psi^{F}_{\alpha}}=e^{-i\epsilon_{\alpha}T}\ket{\psi^{F}_{\alpha}}$, where $\epsilon_{\alpha}\in[0,2\pi]$ are the folded quasi-energies with $\alpha=\{1,2,..,D\}$ and $D=2^{L}$ as the Hilbert-space dimension. To probe localization, we compute the average inverse participation ratio (IPR) $I_{F}=\frac{1}{D}\sum_{\alpha=1}^{D}\sum_{p=1}^{D}|\braket{p}{\psi_{\alpha}^{F}}|^{4}$, where $|p\rangle$ denote product states in the computational basis. The result of $I_{F}$ in Fig. 1(a) shows the localization phase diagram of Floquet eigenstates in the $K$-$J$ parameter space. For weak (no) kick and strong interaction, the Floquet system is in the delocalization phase with vanishing $I_{F}$ when $L\rightarrow\infty$, which corresponds to thermalization and loss of quantum information of initial states. As $K$ is increased, it can inter the MBDL (dynamical localization) phase with finite $I_{F}$ in the interacting (noninteracting) case. In the localization regime, a proper kick strength can induce the DD of certain spins with persistent Rabi oscillation, such as two edge spins for $L=11$ and $K/\Omega=16\pi/25\thickapprox 2.01$ in Fig. 1(a), which will be discussed later. Here the local DD leads to the decrease of $I_{F}$ (and $O_{F}$) with respect to $K$, as depicted in Fig. 1(c).
@@
Figure 1: (Color online) Localization phase diagrams on the $K$-$J$ parameter space obtained by (a) the IPR $I_{F}$ and (b) the staggered magnetization $O_{F}$ averaging over all Floquet eigenstates of $L=11$. The dashed line in (a) corresponding to the values of $I_{F}\approx 0.009$ is plotted to guide the eye and as a rough estimation of the boundary between the delocalization and MBDL phase regions. A proper kick gives rise to the DD of certain spins. The dashed lines with triangles and circles for critical points in (b) represent the phase boundaries determined by finite-size scaling of $O_{F}$ and half-chain entanglement $S_{F}$ of Floquet eigenstates, respectively. (c) $I_{F}$ and $O_{F}$ as a function of $K$ near the DD. (d) $S_{F}$ versus $L$ for various $K$. (e) $\bar{r}$ as a function of $K$ for various $L$. Other parameter is $J=1$ in (c) and $J=2$ in (d,e).
Another quantity to reveal localization in the kicked spin chain is the infinite-time staggered magnetization evolving from an initial Néel state $|\ket{\psi(0)}=|\uparrow\downarrow\uparrow\downarrow...\rangle$, which can be calculated via the Floquet ensemble {|$\{\ket{\psi_{\alpha}^{F}}\}$ as $O_{F}=\frac{1}{L}\sum_{j=1}^{L}\sum_{\alpha=1}^{D}(-1)^{j}|c_{\alpha}|^{2}\bra{\psi^{F}_{\alpha}}\sigma^{z}_{j}\ket{\psi^{F}_{\alpha}}$ Rigol et al. (2008), with the coefficients $c_{\alpha}=\braket{\psi^{F}_{\alpha}}{\psi(0)}$. Figure 1(b) shows the result of $O_{F}$ consistent with the phase diagram in Fig. 1(a). The phase boundary can be estimated by the finite-size scaling of $O_{F}$ (see Fig. 5 in the Supplementary Material (SM) SM ). We also compute the average half-chain entanglement entropy: $S_{F}=-\frac{1}{D}\sum_{\alpha=1}^{D}\operatorname{Tr}\left[\rho^{F}_{\alpha}\log\rho^{F}_{\alpha}\right]$, where $\rho^{F}_{\alpha}=\operatorname{Tr}_{j\leqslant L/2}[\ket{\psi^{F}_{\alpha}}\bra{\psi^{F}_{\alpha}}]$ with even $L$. The finite-size scaling of $S_{F}$ Tang et al. (2021); SM gives roughly consistent boundary between the delocalization and MBDL phases in Fig. 1(b). Figure 1(d) shows the area-law entanglement in the deep MBDL phase since only the area of the bipartitioning contributes to the entanglement Page (1993); Luitz et al. (2015); Kjäll et al. (2014); Pal and Huse (2010), which is different from the volume law in the delocalization phase.
@@
Figure 1(e) indicates the crossover from delocalization to the MBDL by the mean gap-ratio parameter $\bar{r}$ D’Alessio and Rigol (2014); Zhang et al. (2020). The level-spacing statistics of Floquet eigenstates is characterized by $r_{\alpha}=\min\{\delta_{\alpha},\delta_{\alpha+1}\}/\max\{\delta_{\alpha},\delta_{\alpha+1}\}$, where $\delta_{\alpha}=\epsilon_{\alpha+1}-\epsilon_{\alpha}$ is the quasi-energy gap. We obtain $\bar{r}$ up to $L=14$ by averaging $r_{\alpha}$ over all Floquet eigenstates and over 20 realizations of spin chains with a center offset $j_{0}\rightarrow j_{0}+j_{\text{offset}}$ (to lift the level-spacing degeneracy), with $j_{\text{offset}}$ uniformly distributed in the small region $[-0.02,0.02]$. The MBDL phase is indicated by a Poissonian level-spacing distribution with $\bar{r}\approx 0.386$, whereas presents a Wigner-Dyson distribution with $\bar{r}\approx 0.53$ in the delocalized phase. It remains an important task to capture the delocalization-MBDL transition and criticality in the thermodynamics limit Rispoli et al. (2019). However, our results shown in Fig. 1 pinpoint the large parameter region for the MBDL phase in the kicked spin chain.
@@
Dynamical observables.— The (de)localization can be observed via the stroboscopic dynamics of the kicked spin chain. For an initial product state in typical experiments, a direct observation is the spin-up (excitation) density distribution $P_{\uparrow}(j,t)=\left\langle\psi(t)\left|\frac{\sigma^{z}_{j}+\mathbb{I}}{2}\right|\psi(t)\right\rangle$ at the stroboscopic time $t=nT$, where $\mathbb{I}$ is the $2\times 2$ identity matrix. The results of $P_{\uparrow}(j,t)$ for different initial states show similar dynamics SM : The spin chain tends to thermalize in the static case. Turning on periodical kicks up to the MBDL regime, the quantum coherence of initial states preserves without thermalization in the long-time evolution. This disorder-free non-thermalized dynamical phase is due to the kinetic constriction by the quadratic kick potential. The MBDL can be broken down and the thermalization dynamics dominates as $J$ is increased.
@@
Figure 2: (Color online) (a) Time evolution of $\mathcal{I}$ for typical parameters. (b) $\overline{\mathcal{I}}$ as a function of $J$ for $L=13,15,19$. (c) Phase diagram extracted by $\overline{\mathcal{I}}$. (d) Time evolution of subsystem entanglement $S_{A=L/2}$ for various parameters. (e) Saturation entanglement $\bar{S}$ versus $L$. Other parameters are $L=13$ in (a,e) and $K=0.5$ in (b,d,e), and the initial Néel state is $|\ket{\psi(0)}=|\uparrow\downarrow\uparrow\downarrow...\rangle$ in (a-e).
To better characterize the MBDL and extract the localization phase diagram, one can prepare an initial Néel state and measure the evolution of the generalized imbalance $\mathcal{I}(t)=\frac{1}{L}\sum_{j}(-1)^{j+1}P_{\uparrow}(j,t)$. Typical results of $\mathcal{I}(t)$ are shown in Fig. 2(a). In the MBDL (delocalization) phase when $L\rightarrow\infty$, $\mathcal{I}$ remains finite (becomes vanishing) after long time evolution, which is expected to tend to the value of $O_{F}$ in the infinite-time limit. We can use the long-time-averaged imbalance $\overline{\mathcal{I}}$ (averaging over the time interval from $t=100T$ to $200T$ in simulations) to probe (de)localization and extract the phase diagram, as shown in Figs. 2(b) and 2(c), respectively.
@@
Another observable is the dynamics of subsystem entanglement entropy $S_{A}(t)=-\operatorname{Tr}\left[\rho_{A}(t)\log\rho_{A}(t)\right]$ Kaufman et al. (2016); Rispoli et al. (2019); Dong et al. (a); Choi et al. (2023), where $\rho_{A}(t)=\operatorname{Tr}_{j\leq A}[\ket{\psi(t)}\bra{\psi(t)}]$ is the reduced density matrix of subsystem-A. By performing tomography, one can measure $\rho_{A}$ after $n$ kicks to obtain $S_{A}(t)$. Figure 2(d) shows distinct saturation behaviors of the half-chain entanglement $S_{A=L/2}$ for various $L$ in the MBDL and delocalization phases. In finite localized (delocalized) systems, $S_{A=L/2}(t)$ approaches to saturate at late times to certain value independent (linearly dependent) of $L$ Fagotti and Calabrese (2008); Bardarson et al. (2012); Serbyn et al. (2013b); Hart et al. (2021); Lim et al. (2024). The system-size dependence of the saturation entanglement $\bar{S}$, numerically obtained by averaging $S_{A=L/2}(t)$ over the time interval from $t=400T$ to $600T$, is shown in Fig. 2(e). The result confirms the area- and volume-law of the half-chain entanglement in the MBDL and delocalization phases, respectively.
@@
MBDL-protected quantum information.— In the deep MBDL phase, arbitrary states in the energy spectrum are non-thermalized. Due to this strong breaking of ergodicity, the quantum information can be robustly encoded with initial MBDL states. For a generic initial state $|\psi(0)\rangle$, its effective inverse temperature $\beta_{\text{eff}}$ can be obtained by solving $\operatorname{Tr}[H_{F}(\ket{\psi(0)}\bra{\psi(0)}-\rho_{\beta})]=0$, where $\rho_{\beta}=e^{-\beta_{\text{eff}}H_{F}}/\operatorname{Tr}(e^{-\beta_{\text{eff}}H_{F}})$ is the thermal density matrix Bañuls et al. (2011); Chen et al. (2021). We plot $\beta_{\text{eff}}$ as a function of quasi-energies of all product states $\epsilon_{p}=\bra{\psi(0)}H_{F}\ket{\psi(0)}$ for the system of $L=11$ in the MBDL phase and local DD case in Fig. 3(a) and 3(b), respectively. To show the protection of quantum information at high temperatures, we can consider the initial product states with $\beta_{\text{eff}}\approx 0$ near the center of the energy spectra. The non-thermalization dynamics of initial MBDL states can be observed from the long-time coherent evolution of spin density distributions, with an example shown in Fig. 3(d) (and the SM SM ). We can study additional information on the coherent dynamics beyond local observables. We compute the time evolution of fidelity $F_{A}(t)=\left[\operatorname{Tr}\sqrt{\sqrt{\rho_{A}(t)}\rho_{A}(0)\sqrt{\rho_{A}(t)}}\right]^{2}$ and entanglement entropy $S_{A}(t)$ of subsystem-A with $A=4$ for the high-temperature initial states in Figs. 3(e) and 3(f). The subsystem fidelity with high values and low entanglement entropy persists in the long-time evolution.

### v3

@@ Abstract
Dynamically localized states in quantum many-body systems are fundamentally important in understanding quantum thermalization and have applications in quantum information processing. Here we explore many-body dynamical localization (MBDL) without disorders in a non-integrable quantum XY spin chain under periodical and quadratic kicks. We obtain the localization phase regimes with the MBDL and delocalized states and show dynamical observables to extract the phase regimes. For proper kick strengths in the MBDL phase, we reveal a local dynamical decoupling effect for persistent Rabi oscillation of certain spins. Furthermore, we propose the MBDL-protected quantum information at high temperatures, and present an analysis of the dynamical decoupling to obtain the required system parameters for quantum storage. Compared to other non-thermalized states, the disorder-free MBDL states require much fewer repetitions and resources, providing a promising way to protect and store quantum information robust against thermal noises.
Introduction.— Coherent and controllable many-body quantum dynamics plays an important role in quantum matter engineering and quantum information processing. As a consequence of thermalization, however, coherence in generic many-body quantum systems is fragile once they are excited to a finite energy density above ground states Rigol et al. (2008); D’Alessio et al. (2016). There are several ways to avoid thermalization, apart from integrability. One is the many-body localized (MBL) states induced by sufficiently strong disorders Anderson (1958); Nandkishore and Huse (2015); Abanin et al. (2019); Sierant et al. (2025), and the other is quantum many-body scars Bernien et al. (2017); Turner et al. (2018); Bluvstein et al. (2021); Serbyn et al. (2021); Zhang et al. (2022a), which correspond to the strong and weak breaking of ergodicity, respectively. In MBL systems, all excitations are localized and thermalization is strongly suppressed, which can protect coherence of any initial states. The scar states also have a long coherence time as their dynamics are effectively restricted in a subspace. These localized states may provide resources for quantum information processing Huse et al. (2013); Serbyn et al. (2013a, 2014); Bahri et al. (2015); Yao et al. ; Zhang et al. (2018); Dong et al. (a); Naik et al. (2019); Mi and et al. (2022); Schmid et al. (2024); Yarloo et al. (2024).
@@
In this work, we investigate the disorder-free MBDL in a non-integrable kicked XY spin chain and its application in protecting quantum information. For the quantum XY chain with a longitudinal field under periodic transverse-field kicks, we obtain the phase diagram with the MBDL and delocalized phases and show dynamical observables to extract the localization phase regimes. We uncover a local dynamical decoupling (DD) effect under proper kick strengths in the MBDL regime, which enables persistent Rabi oscillation of certain spins. Moreover, we demonstrate the MBDL-protected quantum information at high temperatures, and present an analysis of the local DD to obtain the system parameters for robust quantum storage. Our proposal is realizable with superconducting and Rydberg atom quantum simulators.
@@
where $H_{\text{XY}}=\sum_{j=1}^{L}J\left(\sigma_{j}^{+}\sigma_{j+1}^{-}+\sigma_{j+1}^{+}\sigma_{j}^{-}\right)+\Omega\sigma_{j}^{x}$ is the static Hamiltonian with XY interaction between nearest-neighbour spins and a longitudinal field, $K$ and $T$ denote the strength and period of the transverse-field kicks with a quadratic potential centered at $j_{0}$, respectively. Here $\sigma_{j}^{\pm}=\sigma_{j}^{x}\pm i\sigma_{j}^{y}$ with $\sigma_{j}^{x,y,z}$ being the Pauli matrices of $j$-th spin, $J$ is the interaction strength, and $\Omega$ is the strength of the longitudinal field, which leads to coherent spin flips and breaks $U(1)$ symmetry. Notably, the quadratic form of the periodic kick introduces dynamical constraints on the spin chain, which tend to dynamically localize spins and compete with the XY interaction term. Other kinds of kick potentials, such as a linear potential, can also be chosen to induce the dynamical localization. We will emphasize in protection of quantum coherence raising from $\Omega$ via the MBDL in this kicked XY chain, which was not considered in kicked Ising chains Prosen (2002); Bertini et al. (2018, 2019); Khasseh et al. (2023); Sharma and Bhosale (2024). Hereafter we consider the high-frequency driving by setting $T=1/16$ and $\Omega=\hbar=1$ as the energy (frequency) unit.
@@
where $H_{F}=\frac{i}{T}\ln U_{F}$ is the effective Floquet Hamiltonian. The wave function at time $t=nT$ after $n$ kick cycles is |$\ket{\psi(t)}=U_{F}^{n}\ket{\psi(0)}$. As the model is non-integrable and $U(1)$-symmetry broken, we study the static and dynamical properties using the exact diagonalization Weinberg and Bukov (2019) for $L\leq 14$ and use the time-dependent variational principle (TDVP) Fishman et al. (2022); SM to study the dynamics of large systems up to $L=61$. Without loss of generality, we consider centrosymmetric spin chains by setting $j_{0}=(L+1)/2$ under open boundary condition.
@@
Figure 1: (Color online) Localization phase regimes on the $K$-$J$ parameter space obtained by (a) the averaged half-chain entanglement entropy $S_{F}$ of $L=12$, (b) the staggered magnetization $O_{F}$ and (c) the IPR $I_{F}$ averaging over all Floquet eigenstates of $L=11$. The dashed lines with squares and triangles for critical points in (a) and (b) represent the phase boundaries estimated by the finite-size scaling of $S_{F}$ and $O_{F}$ of Floquet eigenstates SM , respectively. A proper kick gives rise to the DD of certain spins. (d) $S_{F}$ versus $L$ for various $K$. (e) $\bar{r}$ as a function of $K$ for various $L$. Other parameter is $J=1$ in (c) and $J=2$ in (d,e).
Localization phase regimes.— We first study localization properties of many-body Floquet eigenstates {|$\{\ket{\psi^{F}_{\alpha}}\}$. They can be numerically obtained from the eigen equation $U_{F}\ket{\psi^{F}_{\alpha}}=e^{-i\epsilon_{\alpha}T}\ket{\psi^{F}_{\alpha}}$, where $\epsilon_{\alpha}\in[0,2\pi]$ are the folded quasi-energies with $\alpha=\{1,2,..,D\}$ and $D=2^{L}$ as the Hilbert-space dimension. To probe localization phases, we first compute the average half-chain entanglement entropy: $S_{F}=-\frac{1}{D}\sum_{\alpha=1}^{D}\operatorname{Tr}\left[\rho^{F}_{\alpha}\log\rho^{F}_{\alpha}\right]$, where $\rho^{F}_{\alpha}=\operatorname{Tr}_{j\leqslant L/2}[\ket{\psi^{F}_{\alpha}}\bra{\psi^{F}_{\alpha}}]$ with even $L$. The result of $S_{F}$ in Fig. 1(a) shows the localization phase regime in the $K$-$J$ parameter space. For weak (no) kick and strong interaction, the kicked spin chain is in the delocalization phase with large $S_{F}$ obeying the volume law, which corresponds to thermalization and loss of quantum information of initial states. As $K$ is increased, it can inter the MBDL (dynamical localization) phase with small $S_{F}$ obeying the area law in the interacting (noninteracting) case. Another quantity to reveal localization is the infinite-time staggered magnetization evolving from an initial Néel state $|\ket{\psi(0)}=|\uparrow\downarrow\uparrow\downarrow...\rangle$, which can be calculated via the Floquet ensemble {|$\{\ket{\psi_{\alpha}^{F}}\}$ as $O_{F}=\frac{1}{L}\sum_{j=1}^{L}\sum_{\alpha=1}^{D}(-1)^{j}|c_{\alpha}|^{2}\bra{\psi^{F}_{\alpha}}\sigma^{z}_{j}\ket{\psi^{F}_{\alpha}}$ Rigol et al. (2008), with the coefficients $c_{\alpha}=\braket{\psi^{F}_{\alpha}}{\psi(0)}$. Figure 1(b) shows the numerical result of $O_{F}$ in the $K$-$J$ plane, with similar phase regimes shown in Fig. 1(a). We also compute the average inverse participation ratio (IPR) $I_{F}=\frac{1}{D}\sum_{\alpha=1}^{D}\sum_{p=1}^{D}|\braket{p}{\psi_{\alpha}^{F}}|^{4}$, where $|p\rangle$ denote product states in the computational basis. The numerical result in Fig. 1(c) shows small and large $I_{F}$ in the delocalized and MBDL phases, respectively, which agrees with that in Fig. 1(b).
@@
In the localization regime, a proper kick strength can induce the DD of certain spins with persistent Rabi oscillation, such as two edge spins for $L=11$ and $K/\Omega=16\pi/25\thickapprox 2.01$ in Figs. 1(b) and 1(c). Here the local DD leads to the decrease of $O_{F}$ and $I_{F}$ with respect to $K$, and will be discussed later. Fig. 1(d) further shows the area-law entanglement in the deep MBDL phase since only the area of the bipartitioning contributes to the entanglement Page (1993); Luitz et al. (2015); Kjäll et al. (2014); Pal and Huse (2010), which is different from the volume law in the delocalized phase. Figure 1(e) indicates the crossover from delocalization to the MBDL by the mean gap-ratio parameter $\bar{r}$ D’Alessio and Rigol (2014); Zhang et al. (2020). The level-spacing statistics of Floquet eigenstates is characterized by $r_{\alpha}=\min\{\delta_{\alpha},\delta_{\alpha+1}\}/\max\{\delta_{\alpha},\delta_{\alpha+1}\}$, where $\delta_{\alpha}=\epsilon_{\alpha+1}-\epsilon_{\alpha}$ is the quasi-energy gap. We obtain $\bar{r}$ up to $L=14$ by averaging $r_{\alpha}$ over all Floquet eigenstates and over 20 realizations of spin chains with a center offset $j_{0}\rightarrow j_{0}+j_{\text{offset}}$ (to lift the level-spacing degeneracy), with $j_{\text{offset}}$ uniformly distributed in the small region $[-0.02,0.02]$. The MBDL phase is indicated by a Poissonian level-spacing distribution with $\bar{r}\approx 0.386$, whereas presents a Wigner-Dyson distribution with $\bar{r}\approx 0.53$ in the delocalized phase.
Similar to the MBL transition Nandkishore and Huse (2015); Abanin et al. (2019); Sierant et al. (2025), the transition between the MBDL and delocalized (thermal) phases is a dynamical phase transition. The understanding of the transition is still notoriously incomplete due to the lack of analytical treatments and the inherent limitations in numerical simulations. Here we focus on the phase regimes for the MBDL and estimate the phase boundaries by the finite-size scaling of $S_{F}$ and $O_{F}$ for small system sizes with $L\leq 14$ by the exact diagonalization (see the Supplementary Material (SM) SM ). The numerical results give roughly consistent phase boundaries in Figs. 1(a) and 1(b). It remains an important task to capture the criticality near the MBDL-delocalization transition Rispoli et al. (2019). However, our results pinpoint large parameter regions for the MBDL phase in the kicked spin chain.
@@
Dynamical observables.— The (de)localization can be observed via the stroboscopic dynamics of the kicked spin chain. For an initial product state in experiments, a direct observation is the spin-up (excitation) density distribution $P_{\uparrow}(j,t)=\left\langle\psi(t)\left|\frac{\sigma^{z}_{j}+\mathbb{I}}{2}\right|\psi(t)\right\rangle$ at the stroboscopic time $t=nT$, where $\mathbb{I}$ is the $2\times 2$ identity matrix. The results of $P_{\uparrow}(j,t)$ for different initial states show similar dynamics SM : The spin chain tends to thermalize in the static case. Turning on periodical kicks up to the MBDL regime, the quantum coherence of initial states preserves after the long-time evolution. This disorder-free non-thermalized dynamical phase is due to the kinetic constriction by the quadratic kick potential. The MBDL can be broken down and the thermalization dynamics dominates as $J$ is increased.
@@
Figure 2: (Color online) (a) Time evolution of subsystem entanglement $S_{A=L/2}$ for typical parameters. (b) Finite-size scaling of $\bar{S}$. (c) Time evolution of $\mathcal{I}$ for typical parameters. (d) Finite-size scaling of $\overline{\mathcal{I}}$. Insets in (b) and (d) show $\bar{S}$ and $\overline{\mathcal{I}}$ versus $K$ for various $L$, respectively. (e,f) Localization phase regimes with rough boundaries extracted by $\bar{S}$ and $\overline{\mathcal{I}}$, respectively. Other parameters are $L=12$ in (e), $L=11$ in (f) and $J/\Omega=0.5$ in (b,d), and the initial Néel state is $|\ket{\psi(0)}=|\uparrow\downarrow\uparrow\downarrow...\rangle$ in (a-h). The numerical results for $L>14$ are obtained by the TDVP in the localized regime where the entanglement is low Fishman et al. (2022); SM .
@@
To further indicate the MBDL and extract the localization phase regimes in Figs. 1(a) and 1(b), we propose to measure the dynamical observables corresponding to $S_{F}$ and $O_{F}$. Firstly, one can prepare an initial Néel state and measure the evolution of subsystem entanglement entropy $S_{A}(t)=-\operatorname{Tr}\left[\rho_{A}(t)\log\rho_{A}(t)\right]$ Kaufman et al. (2016); Rispoli et al. (2019); Dong et al. (a); Choi et al. (2023), where $\rho_{A}(t)=\operatorname{Tr}_{j\leq A}[\ket{\psi(t)}\bra{\psi(t)}]$ is the reduced density matrix of subsystem-A. By performing tomography, one can measure $\rho_{A}$ after $n$ kicks to obtain $S_{A}(t)$. Figure 2(a) shows distinct saturation behaviors of the half-chain entanglement $S_{A=L/2}$ for various $L$ in the MBDL and delocalized phases. In finite localized (delocalized) systems, $S_{A=L/2}(t)$ approaches to saturate at late times to values independent (linearly dependent) of $L$ Fagotti and Calabrese (2008); Bardarson et al. (2012); Serbyn et al. (2013b); Hart et al. (2021); Lim et al. (2024). Moreover, we can use the long-time-averaged entanglement entropy $\overline{S}$ (averaging over the time interval from $t=400T$ to $600T$ in simulations) to indicate the MBDL, similar as that in Ref. Verdel et al. (2023). The numerical results of $\overline{S}$ versus $K$ for $J=0.5$ and various $L$ are shown in Fig. 2(b). A finite-size scaling analysis of $\overline{S}$ yields the estimated critical point at $K_{c}=0.18$ and critical exponent $\nu=0.65$ for the considered parameters.
Another observable is the generalized imbalance $\mathcal{I}(t)=\frac{1}{L}\sum_{j}(-1)^{j+1}P_{\uparrow}(j,t)$, which approaches to the value of $O_{F}$ in the infinite-time limit. The results of $\mathcal{I}(t)$ for finite chains are shown in Fig. 2(c). In the MBDL (delocalized) phase when $L\rightarrow\infty$, $\mathcal{I}$ remains finite (becomes vanishing) after long time evolution. The long-time-averaged imbalance $\overline{\mathcal{I}}$ (averaging over the time interval from $t=100T$ to $200T$ in simulations) can also be used to detect the MBDL and estimate the critical point. As shown in Fig. 2(d), the finite-size scaling of $\overline{\mathcal{I}}$ yields $K_{c}=0.14$ and $\nu=0.70$, which are close to those in Fig. 2(b). Finally, we plot the results of $\overline{S}$ and $\overline{\mathcal{I}}$ in the $K$-$J$ plane in Figs. 2(e) and 2(f), respectively, for the same system sizes as those in Figs. 1(a) and 1(b). By the finite-size scaling of $\overline{S}$ and $\overline{\mathcal{I}}$ for various values of $J$, we obtain the phase boundaries. These results for different static and dynamical quantities give consistent localization phase regimes and rough boundaries.
@@
MBDL-protected quantum information.— In the deep MBDL phase, arbitrary states in the energy spectrum are non-thermalized. Due to this strong breaking of ergodicity, the quantum information can be robustly encoded with initial MBDL states. For a generic initial state $|\psi(0)\rangle$, its effective inverse temperature $\beta_{\text{eff}}$ can be obtained by solving $\operatorname{Tr}[H_{F}(\ket{\psi(0)}\bra{\psi(0)}-\rho_{\beta})]=0$, where $\rho_{\beta}=e^{-\beta_{\text{eff}}H_{F}}/\operatorname{Tr}(e^{-\beta_{\text{eff}}H_{F}})$ is the thermal density matrix Bañuls et al. (2011); Chen et al. (2021). We plot $\beta_{\text{eff}}$ as a function of quasi-energies of all product states $\epsilon_{p}=\bra{\psi(0)}H_{F}\ket{\psi(0)}$ for the system of $L=11$ in the MBDL phase and local DD case in Fig. 3(a) and 3(b), respectively. To show the protection of quantum information at high temperatures, we can consider the initial product states with $\beta_{\text{eff}}\approx 0$ near the center of the energy spectra. The non-thermalization dynamics of initial MBDL states can be observed from the long-time coherent evolution of spin density distributions, with an example shown in Fig. 3(d) (and the SM SM ). We can study additional information on the coherent dynamics beyond local observables. We compute the time evolution of fidelity $F_{A}(t)=\left[\operatorname{Tr}\sqrt{\sqrt{\rho_{A}(t)}\rho_{A}(0)\sqrt{\rho_{A}(t)}}\right]^{2}$ and entanglement entropy $S_{A}(t)$ of subsystem-A with $A=4$ for the high-temperature initial states in Figs. 3(c) and 3(e). The subsystem fidelity with high values and low entanglement entropy persists in the long-time evolution. Note that some spins in the center of the chain could be non-thermalized and loss local information. To improve the fidelity of initial states, one can increase the ratio between the kick and interaction strengths $K/J$ (see Fig. S8 in the SM SM ).